Two trains going in opposite directions leave at the same time. Train B travels 30 mph faster than train A. In 5 hours the trains are 550 miles apart. Find the speed of each.
Answer by ikleyn(52803)

.
Two trains going in opposite directions leave at the same time.
Train B travels 30 mph faster than train A. In 5 hours the trains are 550 miles apart.
Find the speed of each.
~~~~~~~~~~~~~~~~~~~~~


Let x = the speed of the slower train, in miles per hour.

Then the speed of the faster train is (x+30) mph.


Write the total distance equation

    5x + 5(x+30) = 550  miles.


Simplify and solve it

    5x + 5x = 550 - 150

      10x   =   400

        x   = 400/10 = 40.


ANSWER.  40 mph for the slower train and 40+30 = 70 mph for the faster train.


CHECK.   5*40 + 5*70 = 200 + 350 = 550 miles.   ! Correct !

Solved.

---------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.